OpenStudy (anonymous):

Derivative of f(x) = ((2-3x^2 )^5)/5x

OpenStudy (anonymous):

$$-6 x^2 \left(2-3 x^2\right)^4+\frac{1}{5} \left(2-3 x^2\right)^5$$